Use the discriminant to determine how many solutions there are to the equation.
-0.2xsquared+12x+11=0 Show your work.
Find the solution(s) to the equation. Is your answer in (c) correct?

Answer by stanbon(75887)

Use the discriminant to determine how many solutions there are to the equation.
-0.2xsquared+12x+11=0 Show your work.
-----
b^2-4ac = 12^2-4*-0.2*11 = 152.8
Since the discriminant is > 0, the equation has two unequal
Real Number solutions:
-------------------------------

Find the solution(s) to the equation. Is your answer in (c) correct?
--
x = [-12 +-sqrt(152.8)]/(-0.4)
---
x = [-12 + 12.36]/(-0.4) or x = [-12 - 12.36]/(-0.4)
=======================================
Cheers,
Stan H.